# Enhancing GANs with Contrastive Learning-Based Multistage Progressive Finetuning SNN and RL-Based External Optimization thanks: *GitHub repository: https://github.com/osamaa-mustafa/ImprovingGANs

### v2

@@ V-A Evaluation Metrics
FID Definition
The FID between the distributions of real images $p_{r}$ and generated images $p_{g}$ is defined as:
$\text{FID}(p_{r},p_{g})=\|\mu_{r}-\mu_{g}\|^{2}+\text{Tr}\left(\Sigma_{r}+\Sigma_{g}-2(\Sigma_{r}\Sigma_{g})^{\frac{1}{2}}\right)$ (7)
Where $\mu_{r}$ and $\Sigma_{r}$ are the mean and covariance of the feature vectors for the real images. $\mu_{g}$ and $\Sigma_{g}$ are the mean and covariance of the feature vectors for the generated images. $\|\mu_{r}-\mu_{g}\|^{2}$ is the squared Euclidean distance between the means. Tr denotes the trace of a matrix.
Calculation
To calculate FID in this work, we follow these steps:
  1. 1.
    For computational feasibility and due to continuous reevaluation during experimentation, we set the number of samples in both real and generated data distributions to calculate FID.
  2. 2.
    We select 5000 random images from real data and we generate 5000 images using trained Generator from the GAN.
  3. 3.
    Then we calculate FID using a single FID [11] repository for all of the experiments reported in this work.
@@
We also selected a non-conventional metric to evaluate the proposed framework. Perceptual Path Length (PPL) is deployed to evaluate the smoothness and continuity of the latent space in Generative Adversarial Networks (GANs). Two points Z1 and Z2 are randomly sampled from the latent space. A linear interpolation is performed between these two points to get intermediate latent vectors. For each interpolated latent vector, the GAN generates an image. The perceptual distance between consecutive images is measured. The path length is the sum of these perceptual distances along the interpolation path.
VI Results and Discussion
@@
Ensuring Consistency in Evaluation Metric
In the previous work, researchers have done some remarkable work and reported their results via FID score as a metric. However, during this study, we identified that there is no consistency found in the FID implementation being utilized to report the work. However, it is important to note that FID is calculated based on feature representations extracted using a feature extractor, and a number of different implementations online use a number of different feature extractors trained on different datasets. While some of the works do identify in their work that they are utilizing a custom feature extractor or any modification to report FID scores. To address this issue and maintain consistency in our reported results, we utilize a single FID implementation, and train all the other architectures that we make comparisons with on a single dataset, and report the FID scores for transparency.
@@ Results
We have trained the following architectures on our dataset to benchmark the performance of our proposed solution. The dataset is an open-source challenge dataset, the FID repository is also open-source for reproducibility.
Different variants of GANs exist in the literature. We trained Deep Convolutional Generative Adversarial Network DCGAN for 200 epochs and same configuration as of our proposed model’s training. In addition to a standard GAN, it incorporates convolutional and convolutional-transpose layers to improve the stability and quality. There is an interesting extension of DCGAN, the improved DCGAN [27] in which the authors propose several modifications such as separate real and mini batches, gradient clipping and more, we trained it for same number of epochs and train configuration respectively. It is the same architecture on top of which we propose our solution. Another important variant of GAN is Spectral Normalization GAN (SN-GAN) [21], where authors propose an approach to achieve training stability in a GAN by applying spectral normalization to the weights. This was also trained with the same train configuration. Two important variants that we trained on the selected dataset are Wasserstein GAN with Gradient Penalty (WGAN-GP) [10] and Least Squares GAN (LSGAN) [20]. In wasserstein GAN, the authors propose an approach in which the discriminator as a critic outputs real-valued scores using Wasserstein distance as a distance measure between real and generated images. They also replace weight clipping with gradient penalty. LSGAN utilizes least square loss. In addition to popular GAN variants, in this study we also train a Denoising Diffusion Probabilistic Model (DDPM) [13] on the selected dataset.
TABLE I: FID scores for our proposed and remaining architectures.
Architecture FID Score
DCGAN 57.452
Improved DCGAN 51.256
SNGAN 290.079
@@
LSGAN 203.525
Denoising Diffusion Probalistic Model 70.007
Our Proposed Solution 44.038
TABLE II: PPL values for our proposed and remaining architectures.
@@
TABLE III: KID scores for our proposed and remaining architectures.
@@
TABLE IV: Precision, recall and F1-scores for our proposed and remaining architectures.
@@ VI-A Results
    The model trained on synthetic data, and tested on data gives a test accuracy of 77.41%. Class-wise accuracies are Benign: 75.0% and Invasive: 99.21%.
@@
As the classifier when trained on synthetic data performs much better than when trained on real data, this validates that the proposed GAN framework is able to generalize well and capture significant patterns in the data. It suggests that the synthetic data produced by the proposed GAN is not only representative of the real data but also provides a better diversity and balance of features compared to the real data, especially when the class-wise accuracy of the synthetic trained classifier on the Invasive class is 99.21%. As in this paper, we are proposing this work, in our future work we will test the scale of generalization on very large datasets and deeper networks.
@@
T-sne visualization is performed on real and data distributions overlapping on top of each other. Blue refers to real data points while Red refers to generated data points. PCA is applied as a pre-processing step to reduce the dimensions of feature space to 50. The visualization illustrates that for the proposed model the synthetic (generated) data is well distributed over the real data distribution thus being well representative of the real data. In the rest of the GAN variants, either the two distributions are at a distance, or in the case of improved DCGAN, the synthetic data distribution is concentrated towards the lower left of the real data distribution.
Grad-CAM Visualization using MFT-SNN
We applied grad cam to some of the histopathology patches by utilizing the MFT-SNN feature extractor to check what areas in the image it most gives importance to during feature extraction for similarity computation. As seen in the images, the high-intensity warmer areas are near cells, while other background areas are relatively cooler.
@@
Figure 17: Grad cam visualization - III
Discussion
It can be seen in the Table I, the FID scores for different architectures are reported. The proposed solution outperforms other architectures in the table with a low FID score of 44.038. Although DCGAN has a relatively low FID score, still there are issues with training balance and mode collapse. Other architectures such as SNGAN, LSGAN or WGAN-GP have high FID scores. The diffusion model also has a comparatively low FID score but there are other issues associated with diffusion models such as large inference time and large training data size requirement. Similarly the KID and Precision, Recall, and F1 metrics also report similar results referred to in Tables III & IV. In Table II, the perceptual path length of our proposed model is the lowest of all the corresponding variants reported. A low perceptual path length refers to two qualities:
Qualitative Evaluation
For qualitative evaluation, we developed a Graphical User Interface (GUI). We deployed this GUI to extract K similar images from a selected data distribution for a selected image I. At the backend, we utilized our MFT-SNN to calculate similarity. It is similar to BP-Test [2]. We conducted several different tests thus greater than 10, to check if the proposed model is generating duplicates. For a single test, we noticed a single noisy match being replicated, which did not appear for a model trained on a different dataset. Overall, the qualitative tests proved to be successful with no duplicates and extreme similarities detected in the synthetic data distribution.
VII Conclusion
This study represents a significant contribution to the field of Generative AI, particularly in the context of GANs applied to histopathology image generation. We introduce a novel framework that incorporates Contrastive Learning-based Multistage Progressive Finetuning SNN and RL-based External Optimization. Our findings demonstrate that integrating an external guide during adversarial training enhances the discriminator’s role as a critic, which strengthens the overall adversarial process. This improvement enables the generator to produce higher-quality images while maintaining balanced training to mitigate mode collapse in GANs.
By addressing some inherent issues within GANs, we lay the groundwork for future research. Our goal is to scale up this model using datasets from diverse domains and varying sizes. We encourage the research community to build upon this work and explore further advancements in this direction.
Acknowledgment
I would like to acknowledge Mr. Muhammad Khizer Ali from Charles Sturt University, Australia, and Ms. Nehal Yasin from the University of Southampton, United Kingdom, for their invaluable support and guidance. I would also like to express my gratitude to Dr. Momina Moetesum, my former supervisor, for teaching me research writing, which has been instrumental in conducting this study.

### v3

@@ 4.1 Evaluation Metrics
FID Calculation
To calculate FID in this work, we follow a consistent protocol for computational feasibility and ongoing reevaluation during experimentation. We set the number of samples in both real and generated data distributions and calculate FID using a single FID implementation [8] for all reported experiments, addressing inconsistencies found in previous studies due to varying feature extractors trained on different datasets.
@@
We also selected a non-conventional metric to evaluate the proposed framework. Perceptual Path Length (PPL) is deployed to evaluate the smoothness and continuity of the latent space in GANs. Two points Z1 and Z2 are randomly sampled from the latent space and linear interpolation is performed between these two points to get intermediate latent vectors. The perceptual distance is the sum of these perceptual distances along the interpolation path.
5 Results and Discussion
@@ Results
We trained the Deep Convolutional Generative Adversarial Network (DCGAN) for 200 epochs, using the same configuration as our proposed model. This variant incorporates convolutional and convolutional-transpose layers to enhance stability and quality. We also trained the improved DCGAN [20], which introduces modifications such as separate real and mini-batches and gradient clipping, using identical training epochs and configuration. Another important variant is the Spectral Normalization GAN (SN-GAN) [15], which achieves training stability by applying spectral normalization to the weights, also using the same training configuration. We further trained the Wasserstein GAN with Gradient Penalty (WGAN-GP) [7], where the discriminator outputs real-valued scores based on Wasserstein distance, replacing weight clipping with gradient penalty, and the Least Squares GAN (LSGAN) [14], which utilizes least squares loss. Additionally, we trained a Denoising Diffusion Probabilistic Model (DDPM) [10] on the same dataset.
Table 1: FID scores for architectures.
Architecture FID
DCGAN 57.45
Improved DCGAN 51.26
SNGAN 290.08
@@
LSGAN 203.53
DDPM 70.01
Our Solution 44.04
Table 2: PPL values for our proposed and remaining architectures.
@@
Table 3: KID scores for our proposed and remaining architectures.
@@
Table 4: Precision, recall and F1-scores for our proposed and remaining architectures.
@@ 5.1 Results
    The model trained on synthetic data, and tested on real data gives a test accuracy of 77.41%. Class-wise accuracies are Benign: 75.0% and Invasive: 99.21%.
@@
As the classifier when trained on synthetic data performs much better than when trained on real data, this validates that the proposed GAN framework is able to generalize well and capture significant patterns in the data. It suggests that the synthetic data produced by the proposed GAN is not only representative of the real data but also provides a better diversity and balance of features compared to the real data, especially when the class-wise accuracy of the synthetic trained classifier on the Invasive class is 99.21%. As in this paper, we are proposing this work, in our future work we will test the scale of generalization on much larger datasets and deeper networks.
@@
T-SNE visualization shows overlapping real (blue) and generated (red) data distributions, with PCA reducing the feature space to 50 dimensions. This illustrates that the synthetic data from our proposed model is well-distributed over the real data, while other GAN variants either have distant distributions or, in the case of improved DCGAN, the synthetic data is concentrated towards the lower left of the real data distribution.
Discussion
It can be seen in the Table 1, the FID scores for different architectures are reported. The proposed solution outperforms other architectures in the table with a low FID score of 44.038. Although DCGAN has a relatively low FID score, still there are issues with training balance and mode collapse. Other architectures such as SNGAN, LSGAN or WGAN-GP have high FID scores. The diffusion model also has a comparatively low FID score but there are other issues associated with diffusion models such as large inference time and large training data size requirement. Similarly the KID and Precision, Recall, and F1 metrics also report similar results referred to in Tables 3 & 4. In Table 2, the perceptual path length of our proposed model is the lowest of all the corresponding variants reported. A low perceptual path length refers to two qualities:
We applied Grad-CAM on histopathology patches using the MFT-SNN feature extractor to identify areas prioritized during similarity computation. High-intensity warm areas appear near cells, while background regions remain cooler.
@@ Discussion
5.1.1 Qualitative Evaluation
For qualitative evaluation, we developed a Graphical User Interface (GUI) to extract K similar images from a selected data distribution for a given image I, using our MFT-SNN to calculate similarity, akin to the BP-Test [2]. After conducting over 10 tests, we identified a single noisy match in one instance, which did not appear with models trained on different datasets. Overall, the qualitative tests were successful, revealing no duplicates or extreme similarities in the synthetic data distribution.
6 Conclusion
This study significantly advances Generative AI by introducing a novel framework for GANs in histopathology image generation, integrating Contrastive Learning-based Multistage Progressive Finetuning SNN and RL-based External Optimization. Our findings demonstrate that incorporating an external guide during adversarial training enhances the discriminator’s critical role, resulting in higher-quality images and improved training balance, effectively mitigating mode collapse.
By addressing inherent GAN issues, we lay the groundwork for future research focused on scaling this model with diverse datasets. We encourage the research community to build on this work for further advancements in the field, particularly regarding adversarial resistance, as our approach enhances core optimization in GANs.